Need help with part B
(1 point A matched pairs experiment is carried out, yielding the following data np-54, 2.2, s8.1 Is there evidence, at an a 0.08 level of significance, to conclude that there is a difference between the two sets of measurements? Carry out an appropriate hypothesis test, filing in the information requested A The value of the standardized test statistic -1.995880531 Note: For the next part, your answer should use interval notation. An answer of the form (-oo,a) s expressed (-infty, a), an answer of the form (b, oo) is expressed (b, infty), and an answer of the form (-00, a) U (b,0o) is expressed (infty, aju(b, infty) B The rejection region for the standardized test statisti(-1.425068443, infty) U (1.425 C. The p-value is 0229733942 D. Your decision for the thypothesis test OA. Do Not Reject H OB. Reject H OG. Do Not Reject Ho D. Reject Ho
Answer #1

The rejection region is the critical region.

Here it is \left ( -\infty, -1.785 \right ) \cup \left ( 1.785, \infty \right ) .

Thanks.